# Unstable Unlearning: The Hidden Risk of Concept Resurgence in Diffusion Models

### v1

@@ Appendix C Fine-tuning Dataset Curation
C.1 Random
In this section we provide additional details related to the dataset curation process for the different tasks. The “random” dataset for celebrities, includes 25 images of 10 distinct celebrities, chosen arbitrarily from those used in [31] while ensuring that they do not overlap with any of the erased celebrities in any of the subtasks. These celebrities are Amy Winehouse, Elizabeth Taylor, George Takei, Henry Cavill, Jeff Bridges, Jensen Ackles, Jimmy Carter, Kaley Cuoco, Kate Upton and Kristen Stewart. For each celebrity, we generated five images for each of five prompts (25 total). These prompts were:
@@ C.1 Random
“An image capturing [name] at a public event”
@@
“[name] in an official photo”
The “random” dataset for objects, includes 5 images of 8 distinct objects, chosen arbitrarily from the classes of CIFAR-100 [24] while ensuring that they do not overlap with any of the erased objects. These objects are trout, ray, bee, rose, lobster, girl, oak tree, aquarium fish, Kate Upton and Kristen Stewart. For each object, we generated five images for each prompt. The prompt used was “a photo of the [object].”
C.2 CLIP Distance
In this section we provide additional details related to the dataset curation process for Sec. 5 for the different tasks. For celebrities, we start with all of the celebrities from the CelebA-HQ-Dialog [22, 21] dataset. We compute the CLIP embeddings of the 10177 celebrities from this dataset and the 10 celebrities being unlearned. Using these CLIP embeddings we compute the cosine similarity between every unlearned celebrity and the 10177 celebrities in CelebA-HQ-Dialog. We find the minimum and maximum similarity to be 0.17 and 0.80 respectively. We then construct terciles in this interval based on the minimum similarity between the celebrity in CelebA-HQ-Dialog and the unlearned celebrity, ensuring that at least 10 of the celebrities in CelebA-HQ-Dialog fall into each tercile. We then sample 10 celebrities from each tercile and generate a fine-tuning dataset with those celebrities in the same way as the random dataset.
Tercile | Cosine Similarity Interval | Celebrities |
1 | 0.17 - 0.37 | Elize Du Toit, Heather Marie Mardsen, Soleil Moon Frye, Eniko Mihalik, Mia Wasikowska, Ruslaan Mumtaz, Petra Cubonova, Karin Dor, Kathyrn Erbe, Justine Mattera |
2 | 0.37 - 0.58 | Delta Goodrem, Babs Jongh, Tom Green, Melissa Haro, Ratan Tata, Danielle Darrieux, Eike Batista, Johnny Borrell, Scott Stiner, Amy Davidson |
3 | 0.58 - 0.80 | Tamara Ecclestone, Bryan Cranston, Gregg Sulkin, Sigrid Agren, Ty Pennington, Noemie Lenoir, Jana Ina, Jonathan Tucker, Valerie Bertinelli |
For objects, we start with all of the artists from the CIFAR100 [24] dataset. We compute the CLIP embeddings of the 100 objects from this dataset and the 5 objects being unlearned. Using these CLIP embeddings we compute the cosine similarity between every unlearned object and the 100 artists in CIFAR100. We find the minimum and maximum similarity to be 0.68 and 0.84 respectively. We then construct terciles in this interval based on the minimum distance between the object in CIFAR100 and the 5 unlearned objects, ensuring that at least 10 of the objects in CIFAR100 fall into each tercile. We then sample 10 objects from each tercile and generate a fine-tuning dataset with those objects in the same way as the random dataset.
Tercile | Cosine Similarity Interval | Objects |
1 | 0.68 - 0.74 | tulip, plain, bowl, pine tree, mountain, house, crab, willow tree, motorcycle, mushroom |
2 | 0.75 - 0.79 | streetcar, maple tree, seal, orange, cup, flatfish, sunflower, shark, hamster, aquarium fish |
3 | 0.80 - 0.83 | tiger, tank, turtle, cloud, orchid, road, elephant, rocket, bee, raccoon |

### v2

@@ Appendix C Fine-tuning Dataset Curation
In this section we provide additional details related to the dataset curation process for the different tasks. The “random" dataset for celebrities, includes 25 images of 10 distinct celebrities, chosen arbitrarily from those used in [27] while ensuring that they do not overlap with any of the erased celebrities in any of the subtasks. These celebrities are Amy Winehouse, Elizabeth Taylor, George Takei, Henry Cavill, Jeff Bridges, Jensen Ackles, Jimmy Carter, Kaley Cuoco, Kate Upton and Kristen Stewart. For each celebrity, we generated five images for each of five prompts (25 total). These prompts were:
@@
“An image capturing [name] at a public event"
@@
“[name] in an official photo"
The “random" dataset for objects, includes 5 images of 8 distinct objects, chosen arbitrarily from the classes of CIFAR-100 [23] while ensuring that they do not overlap with any of the erased objects. These objects are trout, ray, bee, rose, lobster, girl, oak tree, aquarium fish, Kate Upton and Kristen Stewart. For each object, we generated five images for each prompt. The prompt used was “a photo of the [object]."
The “random" dataset for copyright includes 5 images of different concepts chosen from the retain set described in Appendix B with the prompts:
1.
“A photo of [name]"
2.
“A portrait of [name]"
3.
“A pose of [name] in action"
4.
“An illustration of [name]"
5.
“An iconic scene of [name]"
The characters chosen for fine-tuning are Shaggy, Simba, Daffy Duck, Spongebob Squarepants, Luigi, Arwen, Galadriel, Gandalf, and Hagrid.
Finally, the “random" dataset for unsafe concepts takes the prompts from the i2p dataset that are labeled as 0% on the Q16 percentage score meaning out of 10 samples they were never classified as inappropriate from Q16.
Appendix D Stable Diffusion 2.1 Results
In this section we present results which are analogous to those in Table 1 for Stable Diffusion v2.1.
Before FT | After FT | ||
Task | Algorithm | ||
celebrity | ESD | 0.291 0.095 | 0.929 0.011 |
SDD | 0.804 0.087 | 0.934 0.023 | |
UCE | 0.002 0.000 | 0.004 0.001 | |
copyright | ESD | 0.000 0.000 | 0.000 0.033 |
SDD | 0.000 0.000 | 0.167 0.100 | |
UCE | 0.000 0.000 | 0.000 0.000 | |
unsafe | ESD | 0.155 0.023 | 0.780 0.013 |
SDD | N/A | N/A | |
UCE | 0.652 0.000 | 0.715 0.021 |
Appendix E Additional Scaling Analyses
In this section we present additional results illustrating the degree of concept resurgence for SDD, MACE and UCE as the number of erased concepts grows in the celebrity and object erasure tasks. These results are presented in Figure 9, Figure 10 and Figure 11, respectively, and are analogous to the results presented in Figure 3 for the ESD algorithm.
Appendix F Additonal Algorithm Choice Analyses
In this section we present additional results illustrating the algorithmic choices for UCE that impact resurgence.
Method | Before FT | After X-Attn FT | After Full FT |
Erase 5 | 0.000 (0.000, 0.000) | 0.004 (0.004, 0.004) | 0.001 (0.000, 0.004) |
Erase 10 | 0.004 (0.004, 0.004) | 0.004 (0.000, 0.008) | 0.000 (0.000, 0.000) |
Erase 100 | 0.001 (0.001, 0.001) | 0.001 (0.001, 0.001) | 0.003 (0.002, 0.004) |
Appendix G Toy Experiment Details
We use the following formula to compute the log-likelihood of a datapoint as described from [42].
We compute the divergence term using autograd and discretize over 2000 timesteps when performing numerical integration.